Circular No. 6016

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET McNAUGHT-HARTLEY (1994n)

The following precise positions have been reported:

1994 UT α_{2000} δ_{2000} m_{1} Observer
July 7.62939 21^{h}36^{m}11.^{s}52 −41^{o}57′13′′.1  Hartley
7.70578 21 36 10.66 −41 57 42.6  _{”}
8.40087 21 36 03.98 −42 02 32.8 16.5 Helin
8.42917 21 36 03.97 −42 02 42.0  _{”}
8.62804 21 36 01.89 −42 03 57.8  Hartley
8.70443 21 36 01.06 −42 04 25.5  _{”}

M. Hartley (Siding Spring). 1.2-m U.K. Schmidt telescope. Measurer R. H.

McNaught. Rather bad seeing on July 8.628 UT. E. Helin, K. Lawrence, J. Alu and K. Yau (Palomar). 0.46-m Schmidt

telescope. Measurer Alu.

The parabolic orbit below, from seven observations July 5–8, is very indeterminate. It is quite probable that the comet is of short period.

T = 1995 Jan. 6.084 TT  ω = 344.^{o}135
Ω = 28.571 }2000.0
q = 1.94029 AU  i = 17.887
1994 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
July 7 21^{h}36.^{m}28 −41^{o}52.^{′}9 2.032 2.909 142.^{o}9 12.^{o}2 16.2
17 21 33.68 −42 59.3 1.910 2.827 148.6 10.8 15.9
27 21 28.64 −43 57.8 1.809 2.747 151.9 10.0 15.7
Aug. 6 21 21.59 −44 40.0 1.730 2.668 151.9 10.3 15.5
16 21 13.44 −44 57.7 1.674 2.591 148.3 11.8 15.3
26 21 05.47 −44 45.1 1.639 2.517 142.4 14.2 15.1

SUPERNOVA 1994S IN NGC 4495 Precise position obtained by A. Boattini and M. Tombelli at the Piaz-

zano Observatory, Florence, on July 5.80 UT: α = 12^{h}28^{m}53^{s}.00, δ = +29^{o}24′37′′.4 (equinox 1950.0).

Provisional CCD photometry obtained by M. Kidger and P. Hammersley with the 0.82-m IAC-80 Telescope at the Teide Observatory on June 10.92 UT: B = 15.33, V = 15.22, R = 15.02, I = 15.24. Visual magnitude estimate by L. Szentaskó, Veresegyház, Hungary: June 17.86 UT, 14.4:.

1994 July 9 Brian G. Marsden